How many years will the following take? $612 to grow to $13,281 if invested at 14.95%, compounded annually. Round the answer to two decimal places.
Answer #1

Let the number of years be n

Interest rate = r = 14.95%

Present Value PV = $612

Future Value FV = $13281

FV = PV(1+r)n

=> 13281 = 612(1+14.95/100)n

=> ln(13281/612) = n ln(1.1495)

=> n = 22.09 years